OpenStudy (anonymous):

a random number generator selects 3 numbers from 1-10. What's the probability all 3 numbers are even?

OpenStudy (dan815):

the probability to get 1 even is half the chance

OpenStudy (dan815):

to get it twice its 1/2 then 1/2 of that so 1/2*1/2 = 1/4 to get it 3 times its half of that again so 1/4 * 1/2 = 1/8

OpenStudy (jebonna):